Hampton Industries had $60,000 in cash at year-end 2018 and $30,000 in cash at year-end 2019. The firm invested in property, plant, and equipment totaling $190,000 — the majority having a useful life greater than 20 years and falling under the alternative depreciation system. Cash flow from financing activities totaled +$130,000. Round your answers to the nearest dollar, if necessary. What was the cash flow from operating activities? Cash outflow, if any, should be indicated by a minus sign. $ If accruals increased by $35,000, receivables and inventories increased by $190,000, and depreciation and amortization totaled $9,000, what was the firm's net income?
Answer #1

Cash at end of year 2018 = $60000, Cash at end of year 2019 = $30000

Net Cash flow for 2019 = Cash at end of year2019 - Cash at end of year 2018 = 30000 - 60000 = -30000

Cash flow from investing activities = - Investment in Property, Plant and Equipment = -$190000

Cash flow from financing activities = 130000

Net Cash for 2019 = Cash flow from operating activities + Cash flow from investing activities + Cash flow from financing activities

-30000 = Cash flow from operating activities -190000 + 130000

Cash flow from operating activities = -30000 + 190000 - 130000 = $30000

Hence Cash flow from Operating activities = $30000

Considering increase in accruals to be increase in Current liabilities , Increase in accruals = $35000,

Increase in accounts receivable and inventories = $190000,

Depreciation and Amortization = $9000

Cash flow from operating activities = Net Income + Depreciation & Amortization - Increase in accounts receivable and inventories + Increase in accruals

30000 = Net income + 9000 - 190000 + 35000

Net income = 190000 -9000 -35000 + 30000 = 176000

Hence Net income = $176000